# Preferences Evolve And So Should Your Bandits: Bandits with Evolving States for Online Platforms

### v2

@@ 5 Fast State Evolution: $\lambda\in[\widetilde{\Theta}(1-1/\sqrt{T}),1]$
Because $q_{t}=b_{i_{t-1}}$ (i.e., the state at $t$ depends on only the action played at $t-1$), playing a sequence $\{i_{1},i_{2},\dots,i_{N^{\prime}}\}$ repeatedly $M$ times gives the same reward in expectation times $M$. In other words, if $\{\widetilde{q}_{1},\dots,\widetilde{q}_{N^{\prime}}\}$ is the sequence of induced states as a result of playing $\{i_{1},\dots,i_{N^{\prime}}\}$, and $q_{t}^{\prime}$ the sequence of induced states for repeating $\{i_{1},\dots,i_{N}^{\prime}\}$ for $M$ times, we have that:
@@
$\underbrace{ABABA\cdots AB{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}A}-{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}C}CC\cdots C}_{\text{Batch }1}-\underbrace{ABABA\cdots AB{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}A}-{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}C}CC\cdots C}_{\text{Batch }2}-\cdots$ (14)
@@
Input. Set of meta arms $\mathcal{A}$, rounds $U_{\beta}.$ Initialize set of unexplored active arms: $\mathcal{A}^{\prime}\leftarrow\mathcal{A}$. while $\mathcal{A^{\prime}}\neq\emptyset$ do
       Choose a random meta-arm $(i\diamond j)\in\mathcal{A}.$ Change arms by playing arm $i$.   // ‘‘Initialize’’ meta-arm by playing $i$.
@@
2             Play arm $j$ and observe reward. Play arm $i$ and observe reward.
3      Update unexplored active meta-arms: $\mathcal{A^{\prime}}\leftarrow\mathcal{A^{\prime}}\setminus{(i\diamond j)}$. while $\exists(x\diamond i)$ or $(i\diamond x)$ pair $\in\mathcal{A}^{\prime}$ do
@@
5                   Play arm $x$ and observe reward. Play arm $i$ and observe reward.
@@
Pick any $(i\diamond j)\in\mathcal{A}^{\prime}$. Algorithm 3 makes $1$ “switch” for the first pair $(i\diamond j)$. Assume that the last action played was action $i$. Then, to explore the meta-arms that are still in $\mathcal{A}^{\prime}$ and include $i$ the algorithm does not make any “switch” (i.e., does not throw away any samples). This is repeated until there are no more meta-arms in $\mathcal{A}^{\prime}$. There are at most $K(K+1)/2$ meta-arms and at most $K$ different arms in these meta-arms. Thus, Algorithm 3 does at most $K$ switches for each batch $B$. ∎
Input. Number of batches $B=2\log T$, $K$ arms, time horizon $T$. Set $w=T^{1/B}=\sqrt{e}$ and generate $M=K(K+1)/2$ meta-arms $(a_{i}\diamond a_{j})$, with $i\leq j,(i,j)\in[K]^{2}$. Set active meta-arms $\mathcal{A}=\{(i\diamond j)\mid i\leq j,i,j\in[K]\}$.   // $|\mathcal{A}|=M$ initially
1 For $i\leq j\in[K]$ initialize estimated means $\hat{\mu}_{(i\diamond j)}=0$. for batch $\beta=1$ to $B-1$ do
@@ 3             Break
      Play all $(i\diamond j)\in\mathcal{A}$ for $U_{\beta}=\lfloor{w^{\beta}\rfloor}$ times according to Algorithm 3.   // contains $12U_{\beta}+$ 1 actions
4       Drop the first reward observation that has expectation $r_{i}q_{0}$, where $q_{0}=1$. Pair the other observations into $U_{\beta}$ groups of size $2$. Update $\hat{\mu}_{(i\diamond j)}$ using these new $U_{\beta}$ observations (sample mean). Update the number of observations of all existing meta-arms according to $c_{\beta}=\sum_{l=1}^{\beta}U_{l}$. for each active arm $(i\diamond j)$ in $\mathcal{A}$ do
@@ 5.2 Evolution Rate $\lambda\in[\widetilde{\Theta}(1-1/\sqrt{T}),1)$
When the evolution rate is equal to $\lambda=1-\epsilon$, then after playing an arm $I_{t}$ the state becomes almost $I_{t}^{\prime}s$ baseline reward: $q_{t+1}=\epsilon q_{t}+(1-\epsilon)\cdot b_{I_{t}}$. Thus, playing the best meta-arm as defined in the previous subsection in near optimal.
@@ 6.2 Unknown $\lambda$
When $\lambda$ is originally unknown, we find an algorithm (Algorithm 8 in the Appendix) that can guarantee sublinear regret under one of the two assumptions: (A1) $\max_{i,j\in[K]}r_{i}|b_{i}-b_{j}|>\tilde{\omega}(1/T^{-1/3})$ or (A2) $\lambda\notin(\widetilde{\Theta}(1/T),\widetilde{\Theta}(K^{1/3}/T^{1/3})]$. Assumption (A1) is a “discrepancy assumption” that intuitively says that there exist two arms whose ES are more than $1/T^{1/3}$ away.
@@
Under either assumption (A1) or (A2), Algorithm 8 incurs regret $\widetilde{\mathcal{O}}({K^{1/3}T^{2/3}})$.
@@ 7 Discussion
For the second avenue for future research, it is an open question how to obtain better bounds for the agnostic $\lambda$ case. Specifically, we think that a very interesting idea would be to try to obtain optimal regret bounds when you are given an original prediction about how big $\lambda$ is, while being robust to potentially adversarial information. The hardness that one would need to overcome would be to find a way to distinguish between good and bad predictions, without having to resort to our approach that alternates a specific set of arms until we have converged.
@@ Appendix B Appendix for Section 3
Let $S$ be the sequence of arms the Algorithm 1 returns and $\pi^{\star}$ the optimal sequence of arms as usual. First, we will prove that achieves $(1-\epsilon)$ approximation and then that it has time complexity $\mathcal{O}(KT^{2}/\epsilon)$. It holds that $\epsilon\left\lfloor\frac{1}{\epsilon}r_{i}\cdot q\right\rfloor\leq r_{i}\cdot q\leq\epsilon(\left\lfloor\frac{1}{\epsilon}r_{i}\cdot q\right\rfloor+1)$ and $|S|=|\pi^{\star}|=T$. Hence:
$\sum_{i\in S}r_{i}q$ $\geq\epsilon\sum_{i\in S}\left\lfloor\frac{1}{\epsilon}r_{i}\cdot q\right\rfloor\geq\epsilon\sum_{i\in\pi^{\star}}\left\lfloor\frac{1}{\epsilon}r_{i}\cdot q\right\rfloor\geq\sum_{i\in\pi^{\star}}r_{i}q-\epsilon\textsf{OPT}\geq\textsf{OPT}-\epsilon\textsf{OPT}=(1-\epsilon)\textsf{OPT}.$
@@
For the rounds that pass while we are on lines 3 – 17 of Algorithm 2, we pick up regret at most $1$ at each of them. Hence, the regret picked up in total equals to the number of rounds between these lines which are $2c(\lambda)\cdot\log(1/\lambda\varepsilon)\cdot K\cdot M$.
We next define events $\mathcal{E}_{r,i}=\{|\widehat{r}_{i}-r_{i}|\leq\delta\}$ and $\mathcal{E}_{b,i}=\{|\widehat{b}_{i}-b_{i}|\leq\delta\}$ for all $i\in[K]$. Then, conditional on the event =$\mathcal{E}=\{\cap_{i\in[K]}(\mathcal{E}_{r,i}\,\,\text{and}\,\,\mathcal{E}_{v,i})\}$ and due to Lemma 3.3, the regret picked up for all the remaining rounds after feeding estimates $\{(\widehat{r}_{i},\widehat{b}_{i})\}_{i\in[K]}$ to the dynamic programming procedure is at most $\delta T$. As a result, from the law of total probability, the regret for all $T$ rounds:

### v4

@@ 5 Fast State Evolution: $\lambda\in[\widetilde{\Theta}(1-1/\sqrt{T}),1]$
Because $q_{t}=b_{I_{t-1}}$ (i.e., the state at $t$ depends on only the action played at $t-1$), playing a sequence $\{i_{1},i_{2},\dots,i_{N^{\prime}}\}$ repeatedly $M$ times, for $N^{\prime}>0$ gives the same reward in expectation times $M$. In other words, if $\{\widetilde{q}_{1},\dots,\widetilde{q}_{N^{\prime}}\}$ is the sequence of induced states as a result of playing $\{i_{1},\dots,i_{N^{\prime}}\}$, and $q_{t}^{\prime}$ the sequence of induced states for repeating $\{i_{1},\dots,i_{N^{\prime}}\}$ for $M$ times, we have that:
@@
$\underbrace{ABABA\cdots AB{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}A}-{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}C}CC\cdots C}_{\text{Batch }1}-\underbrace{ABABA\cdots AB{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}A}-{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}C}CC\cdots C}_{\text{Batch }2}-\cdots$ (14)
@@
Input. Set of meta arms $\mathcal{A}$, rounds $U_{\beta}.$
Initialize set of unexplored active arms: $\mathcal{A}^{\prime}\leftarrow\mathcal{A}$.
while $\mathcal{A^{\prime}}\neq\emptyset$ do
       Choose a random meta-arm $(i\diamond j)\in\mathcal{A}.$
       Change arms by playing arm $i$.
        // ‘‘Initialize’’ meta-arm by playing $i$.
@@
2             Play arm $j$ and observe reward.
3             Play arm $i$ and observe reward.
4
5      Update unexplored active meta-arms: $\mathcal{A^{\prime}}\leftarrow\mathcal{A^{\prime}}\setminus{(i\diamond j)}$.
6       while $\exists(x\diamond i)$ or $(i\diamond x)$ pair $\in\mathcal{A}^{\prime}$ do
@@
8                   Play arm $x$ and observe reward.
9                   Play arm $i$ and observe reward.
10
@@
Pick any $(i\diamond j)\in\mathcal{A}^{\prime}$. Algorithm 3 makes $1$ “switch” for the first pair $(i\diamond j)$. Assume that the last arm played was arm $i$. Then, to explore the meta-arms that are still in $\mathcal{A}^{\prime}$ and include $i$, the algorithm does not make any “switch” (i.e., does not throw away any samples). This is repeated until there are no more meta-arms in $\mathcal{A}^{\prime}$. There are at most $K(K+1)/2$ meta-arms and at most $K$ different arms in these meta-arms. Thus, Algorithm 3 makes at most $K$ switches for each batch $B$. ∎
Input. Number of batches $B=2\log T$, $K$ arms, time horizon $T$.
Set $w=T^{1/B}=\sqrt{e}$ and generate $M=K(K+1)/2$ meta-arms $(a_{i}\diamond a_{j})$, with $i\leq j,(i,j)\in[K]^{2}$.
Set active meta-arms $\mathcal{A}=\{(i\diamond j)\mid i\leq j,i,j\in[K]\}$.
  // $|\mathcal{A}|=M$ initially
1 For $i\leq j\in[K]$ initialize estimated means $\hat{\mu}_{(i\diamond j)}=0$.
2 for batch $\beta=1$ to $B-1$ do
@@ 4             Break
      Play all $(i\diamond j)\in\mathcal{A}$ for $U_{\beta}=\lfloor{w^{\beta}\rfloor}$ times according to Algorithm 3.
        // contains $12U_{\beta}+$ 1 actions
5       Drop the first reward observation that has expectation $r_{i}q_{0}$, where $q_{0}=1$.
6       Pair the other observations into $U_{\beta}$ groups of size $2$.
7       Update $\hat{\mu}_{(i\diamond j)}$ using these new $U_{\beta}$ observations (sample mean).
8       Update the number of observations of all existing meta-arms according to $c_{\beta}=\sum_{l=1}^{\beta}U_{l}$.
9      for each active arm $(i\diamond j)$ in $\mathcal{A}$ do
@@
As for the lower bound of the regret achievable by any no regret algorithm when $\lambda=1$, we have the following:
Lemma 5.4.
Any algorithm for B-DES with $\lambda=1$ incurs regret at least $\Omega(K\sqrt{T})$.
Proof.
Fix two arms $i^{\star}\in[K]$ and $j^{\star}\in[K]$ such that $i^{\star}\neq j^{\star}$. For parameters $\varepsilon_{1},\varepsilon_{2}\in[0,1]$, we define the following instance $\mathcal{I}(i^{\star},j^{\star},\varepsilon_{1},\varepsilon_{2})$:
$\mathcal{I}(i^{\star},j^{\star},\varepsilon_{1},\varepsilon_{2})=\begin{cases}(r_{i},b_{i})=(\frac{1-\varepsilon_{1}}{2},1-\varepsilon_{2})&\textrm{for all arms}\ i\neq\{i^{\star},j^{\star}\}\\ (r_{i},b_{i})=(\frac{1-\varepsilon_{1}}{2},1)&\textrm{for arm}\ i=j^{\star}\\ (r_{i},b_{i})=(\frac{1}{2},1-\varepsilon_{2})&\textrm{for arm}\ i=i^{\star}\end{cases}$
We are going to prove the lower bound for the above instance. [^6] There is an alternative family of instances that achieves the same lower bound, where $i^{\star}=j^{\star}$ and $(r_{i}^{\star},b_{i}^{\star})=(r_{j}^{\star},b_{j}^{\star})=(1/2,1)$.
We first identify the optimal sequence of arms on instance $\mathcal{I}(i^{\star},j^{\star},\varepsilon_{1},\varepsilon_{2})$. Since $\lambda=1$, from Lemma 5.2 we know that the optimal sequence will be a cycle of at most $2$ arms. Observe that the optimal meta-arm is $(i^{\star}\diamond j^{\star})$. This is because for any $(i\diamond j)$ such that $i\neq\{i^{\star},j^{\star}\}$ the expected reward is: $(1-\varepsilon_{1})(1-\varepsilon_{2})/2$; for any $(i^{\star}\diamond j)$ such that $j\neq\{i^{\star},j^{\star}\}$ the expected reward is: $(1-\varepsilon_{1})(1-\varepsilon_{2})/4+(1-\varepsilon_{2})/4$; for any $(i\diamond j^{\star})$ such that $i\neq\{i^{\star},j^{\star}\}$ the expected reward is: $(1-\varepsilon_{1})/4+(1-\varepsilon_{1})(1-\varepsilon_{2})/4$; for $(j^{\star}\diamond j^{\star})$ the expected reward is: $1-\varepsilon_{1}$; for $(i^{\star}\diamond i^{\star})$ the expected reward is: $1-\varepsilon_{2}$; for $(i^{\star},j^{\star})$ the expected reward is: $(1-\varepsilon_{1})(1-\varepsilon_{2})/4+1/4$.
Constraints imposed by the above (in order for $(i^{\star}\diamond j^{\star})$ to be the opt meta-arm):
$(1-\varepsilon_{1})(1-\varepsilon_{2})\leq 1$ (true for any $\varepsilon_{1},\varepsilon_{2}$)
$(1-\varepsilon_{2})/2\leq 1/2$ (true for any $\varepsilon_{1},\varepsilon_{2}$)
$41-\varepsilon_{1}\leq(1-\varepsilon_{1})(1-\varepsilon_{2})/4+1/$ 4
$41-\varepsilon_{2}\leq(1-\varepsilon_{1})(1-\varepsilon_{2})/4+1/$ 4
Note $(i^{\star},j^{\star})$ is the best pair of arms to played (i.e., best meta-arm). Let us also define the instance $I_{i^{\star},j^{\star},\varepsilon_{1},\varepsilon_{2}}$ where the $\varepsilon_{1},\varepsilon_{2}$ are chosen according to following policy: $\varepsilon_{1}=\varepsilon_{2}=1$ if algorithm $\mathcal{A}$ explores less than $O(K^{2})$ meta-arms, else $\varepsilon_{2}=0$ and $\varepsilon_{1}$ will be adjusted in the analysis. We move forward by fixing the aforementioned instance.
We prove that any algorithm that does not explore all meta-arms has linear regret. Fix an algorithm $\mathcal{A}$ that explores less than $O(K^{2})$ arms. Recall that only the products $r_{i}\cdot b_{j}$ are observable. If the algorithm explores less than $O(K^{2})$ meta-arms then, there is a constant $c>0$, such that the set of arms $S$ that it explores is less than $K^{2}/c$. Also, all the mean values of the meta-arms are equal to 0 (since $\varepsilon_{1}=1,\varepsilon_{2}=1$) except from $(i^{\star}\diamond j^{\star})$, which is $1/2$. Thus, the expected regret of these algorithms is $\mathbb{E}[R_{\textsc{DES}}(T)]\geq\Pr[(j\diamond k)\notin S]\cdot T\geq(1/2-1/2c)\cdot T=\Omega(T)$.
Now we study the case where an algorithm $\mathcal{A}$ explores all meta-arms. The rest of the proof is a variation of the stochastic MAB lower bound proof. As we mentioned before, we choose $\varepsilon_{2}=0$. So, we have $K^{2}-K$ Bernoulli distributions $p_{(i\diamond j)}=\textrm{Bern}((1-\varepsilon_{1})/2)$ and $K$ Bernoulli distribution $p_{(i^{\star}\diamond j^{\star})}=\textrm{Bern}(1/2)$. The Kullback-Leibler divergence of $p_{(i\diamond j)}$, $p_{(i^{\star}\diamond j^{\star})}$ equals:
$KL(p_{(i,j)}||p_{(i^{\star}\diamond j^{\star})})$ $=\sum_{x\in\Omega}p_{(i\diamond j)}(x)\ln\left(\frac{p_{(i\diamond j)}}{p_{(i^{\star}\diamond j^{\star})}}\right)$
$=\left(\frac{1-\varepsilon_{1}}{2}\right)\ln(1-\varepsilon_{1})+\left(\frac{1+\varepsilon_{1}}{2}\right)\ln(1+\varepsilon_{1})$
$=\frac{1}{2}\left(\ln(1+\varepsilon_{1})+\ln(1-\varepsilon_{1})\right)+\frac{\varepsilon_{1}}{2}(\ln(1+\varepsilon_{1})-\ln(1-\varepsilon_{1}))$ ($\ln(1-\varepsilon_{1}^{2})<0$)
$\leq\frac{\varepsilon_{1}}{2}(\ln(1+\varepsilon_{1})-\ln(1-\varepsilon_{1}))$
$=\frac{\varepsilon_{1}}{2}\ln\left(\frac{1+\varepsilon_{1}}{1-\varepsilon_{1}}\right)=\frac{\varepsilon_{1}}{2}\ln\left(1+\frac{2\varepsilon_{1}}{1-\varepsilon_{1}}\right)$ ($\ln(1+x)<x$)
$\leq\frac{\varepsilon_{1}^{2}}{1-\varepsilon_{1}}\leq 2\varepsilon_{1}^{2},{}$ (16)
where the last inequality applies for $\varepsilon_{1}\leq 1/2$. Let $n_{(i\diamond j)}$ be the number of samples of meta-arm $(i\diamond j)$. Also let instance $\mathcal{I}_{1/2}=\{r_{i}=b_{i}=1/2\ \forall i\in[K]\}$. For any algorithm $\mathcal{A}$, after $t$ rounds it holds that:
$\sum_{(i\diamond j)\in K^{2}}\mathbb{E}[n_{(i\diamond j)}|\mathcal{I}_{1/2}]=t.$
Hence, the set $J_{1}=\{(i\diamond j):\mathbb{E}[(i\diamond j)|\mathcal{I}_{1/2}]\leq 2K^{2}/3\}$ has at least $32K^{2}/$ 3 elements and the set $J_{2}=\{(i\diamond j):\Pr[(i\diamond j)\neq(i^{\star}\diamond j^{\star})|\mathcal{I}_{1/2}]\leq 3/K^{3}\}$ has at least $32K^{2}/$ 3 meta-arms. Then, $J=J_{1}\cap J_{3}$ has at least $K^{2}/3$ meta arms. If $(i\diamond j)\in J$ and $100t\leq\frac{K^{2}}{\varepsilon^{2}}$:
$\Pr[(i\diamond j)=(i^{\star}\diamond j^{\star})|\mathcal{I}_{i^{\star},j^{\star},\varepsilon_{1},0}]$ $\leq\Pr[(i\diamond j)=(i^{\star}\diamond j^{\star})|\mathcal{I}_{1/2}]$
$\quad\quad+\left|\Pr[(i\diamond j)=(i^{\star},j^{\star})|\mathcal{I}_{i^{\star},j^{\star},\varepsilon_{1},0}]-\Pr[(i\diamond j)=(i^{\star}\diamond j^{\star})|\mathcal{I}_{1/2}]\right|$
$=\frac{3}{K^{2}}+\sqrt{\frac{1}{2}KL(p_{(i\diamond j)}||p_{(i^{\star}\diamond j^{\star})})}$ (Pinsker’s Inequality)
$\leq\frac{3}{K^{2}}+\frac{1}{2\sqrt{2}}<\frac{1}{2}$ (16)
Thus, in some round $t\leq\frac{K^{2}}{100\varepsilon_{1}^{2}}$: $\mathbb{E}[\mu_{(i\diamond j)}]\leq O(\mu_{(i^{\star}\diamond j^{\star})}-\varepsilon_{1})$, which implies that: $\mathbb{E}[R_{\textsc{DES}}(T)]\geq\Omega(K\sqrt{T})$, for $T\leq 100K^{2}/\varepsilon_{1}^{2}$. ∎
@@ 5.2 Evolution Rate $\lambda\in[\widetilde{\Theta}(1-1/\sqrt{T}),1)$
When the evolution rate is equal to $\lambda=1-\epsilon$, then after playing an arm $I_{t}$ the state becomes almost $b_{I_{t}}$, i.e., $q_{t+1}=\epsilon q_{t}+(1-\epsilon)\cdot b_{I_{t}}$. Thus, playing the best meta-arm as defined in the previous subsection in near optimal.
@@ 6.2 Unknown $\lambda$
When $\lambda$ is originally unknown, we present an algorithm (Algorithm 8 in the Appendix) that can guarantee sub-linear regret for $\lambda$in $[0,\tilde{\Theta}(1/T^{2})]\cup[\tilde{\Theta}(K^{2/3}/T^{2/3}),1]$. For $\lambda=T^{-a/b}$ with $b<a<2b$ and for $\lambda=\tilde{\Theta}(1/T)$ we have $R_{\textsc{DES}}(T)=T^{b/a}$ and $(1-1/e)\textsf{OPT}$ respectively as before. For $\lambda\in(\tilde{\Theta}(1/T),\tilde{\Theta}((K/T)^{2/3}))$, we have linear regret.
@@
For $\lambda\notin(\tilde{\Theta}(1/T^{2}),\tilde{\Theta}((K/T)^{2/3}))$, Algorithm 8 incurs regret $\widetilde{\mathcal{O}}({K^{1/3}T^{2/3}})$.
@@ 7 Discussion
For the second avenue for future research, it is an open question how to obtain better bounds for the agnostic $\lambda$ case. Specifically, we think that a very interesting idea would be to try to obtain optimal regret bounds when you are given an original prediction about how big $\lambda$ is, while being robust to potentially adversarial information. The hardness that one would need to overcome would be to find a way to distinguish between good and bad predictions, without having to resort to our approach which alternates a specific set of arms until we have converged.
@@ Appendix B Appendix for Section 3
Let $\widehat{i}^{\star}_{1},\dots,\widehat{i}^{\star}_{T}$ be the sequence returned by the the DP algorithm, then,
$\textsf{DP}\triangleq\sum_{t\in[T]}\left[(1-\lambda)^{t}+\lambda\sum_{s=1}^{t}(1-\lambda)^{t-s}\cdot b_{\widehat{i}^{\star}_{s}}\right]\cdot r_{\widehat{i}^{\star}_{t}}\neq\widehat{\textsf{DP}}.$
Now we are going to bound the difference |DP^DP|^$-\textsf{DP}|$ DP |:
|DP^DP|^$-\textsf{DP}|$ DP | $=\sum_{t\in[T]}\left[(1-\lambda)^{t}+\lambda\sum_{s=1}^{t}(1-\lambda)^{t-s}\cdot b_{\widehat{i}^{\star}_{s}}\right]\cdot r_{\widehat{i}^{\star}_{t}}$
$\quad\quad\quad\quad\quad-\arg\max_{i_{1},\dots,i_{T}}\sum_{t\in[T]}\left[(1-\lambda)^{t}+\lambda\sum_{s=1}^{t}(1-\lambda)^{t-s}\cdot\hat{b}_{i_{s}}\right]\cdot\hat{r}_{i_{t}}$
$\leq\sum_{t\in[T]}\left[(1-\lambda)^{t}+\lambda\sum_{s=1}^{t}(1-\lambda)^{t-s}\cdot b_{\widehat{i}^{\star}_{s}}\right]\cdot r_{\widehat{i}^{\star}_{t}}$
$\quad\quad\quad\quad\quad-\sum_{t\in[T]}\left[(1-\lambda)^{t}+\lambda\sum_{s=1}^{t}(1-\lambda)^{t-s}\cdot\hat{b}_{\widehat{i}^{\star}_{s}}\right]\cdot\hat{r}_{\widehat{i}^{\star}_{t}}$
$\leq\sum_{t\in[T]}\left[(1-\lambda)^{t}+\lambda\sum_{s=1}^{t}(1-\lambda)^{t-s}\cdot b_{\widehat{i}^{\star}_{s}}\right]\cdot r_{\widehat{i}^{\star}_{t}}$
$\quad\quad\quad\quad\quad-\sum_{t\in[T]}\left[(1-\lambda)^{t}+\lambda\sum_{s=1}^{t}(1-\lambda)^{t-s}\cdot(\hat{b}_{\widehat{i}^{\star}_{s}}-\delta)\right]\cdot(\hat{r}_{\widehat{i}^{\star}_{t}}-\delta)$
$\leq\sum_{t\in[T]}\left[(1-\lambda)^{t}\delta+\lambda\sum_{s=1}^{t}(1-\lambda)^{t-s}\cdot 2\delta\right]$
$\leq 2\delta T.$
Proof of Lemma 3.5.
Let $S=\{I_{t}\}_{t\in[T]}$ be the sequence of arms the Algorithm 1 returns and $\pi^{\star}$ the optimal sequence of arms as usual. First, we will prove that it achieves $(1-\epsilon)$ approximation and then that it has time complexity $\mathcal{O}(KT^{2}/\epsilon)$. For any discretization factor $\epsilon$, it holds that $\epsilon\left\lfloor\frac{1}{\epsilon}r_{i}\cdot q\right\rfloor\leq r_{i}\cdot q\leq\epsilon(\left\lfloor\frac{1}{\epsilon}r_{i}\cdot q\right\rfloor+1)$, for any state $q$. The lengths of the sequences $S$ and ${\pi^{\star}}$ are equal: $|S|=|\pi^{\star}|=T$. Hence:
$\sum_{I_{t}\in S}r_{I_{t}}q_{t}^{S}$ $\geq\epsilon\sum_{I_{t}\in S}\left\lfloor\frac{1}{\epsilon}r_{I_{t}}\cdot q_{t}^{S}\right\rfloor\geq\epsilon\sum_{I_{t}\in\pi^{\star}}\left\lfloor\frac{1}{\epsilon}r_{I_{t}}\cdot q_{t}^{\pi^{\star}}\right\rfloor$
$\geq\sum_{I_{t}\in\pi^{\star}}r_{I_{t}}q_{t}^{\pi^{\star}}-\epsilon\textsf{OPT}\geq\textsf{OPT}-\epsilon\textsf{OPT}=(1-\epsilon)\textsf{OPT},$
where the second inequality is because $S$ is the optimal sequence on the discretized grid and $\pi^{\star}$ is the optimal sequence in the continuous space. So, when comparing the total state-augmented reward in the discretized grid, $S$ is better than $\pi^{\star}$.
@@
For the rounds that pass while we are in lines 3 – 17 of Algorithm 2, we pick up regret at most $1$ at each of them; these are the “explore” rounds. Hence, the regret picked up in the explore rounds is $\textsf{Explr}=2c(\lambda)\cdot\log(1/\lambda\varepsilon)\cdot K\cdot M$, and from the definition of regret:
$R_{\textsc{DES}}(T)\leq\textsf{OPT}-\textsf{Explr}-\sum_{t\in[T]}\left[(1-\lambda)^{t}+\lambda\sum_{s\in[t-1]}(1-\lambda)^{t-1-s}b_{I_{t}}\right]\cdot r_{I_{t}}$
Note that $\{I_{t}\}_{t\in[T]}$ is the sequence of arms chosen by the dynamic programming to which we have fed the estimators $\{(\widehat{r}_{i},\widehat{b}_{i})\}_{i\in[K]}$ computed in the explore phase; we slightly alter notation and write $\{I_{t}\}\triangleq\{\hat{i}_{t}^{\star}\}$ to make it clear that the dynamic programming solution is based on the estimators. To match the notation of Lemmas 3.3 and 3.4, we write: $\textsf{DP}:=\sum_{t\in[T]}\left[(1-\lambda)^{t}+\lambda\sum_{s\in[t-1]}(1-\lambda)^{t-1-s}b_{\hat{i}^{\star}_{t}}\right]\cdot r_{\hat{i}^{\star}_{t}}$ and hence:
$R_{\textsc{DES}}(T)\leq\textsf{OPT}-\textsf{Explr}-\textsf{DP}$
Let $\widehat{\textsf{DP}}\triangleq\max_{i_{1},\dots,i_{T}}\sum_{t\in[T]}\left[(1-\lambda)^{t}+\lambda\sum_{s\in[t-1]}(1-\lambda)^{t-1-s}\widehat{b}_{i_{t}}\right]\cdot\widehat{r}_{i_{t}}$, i.e., $\widehat{\textsf{DP}}$ corresponds to the optimal value of the dynamic programming algorithm for the estimators. Note that the sequence of arms that maximizes $\widehat{\textsf{DP}}$ is $\{\widehat{i}_{1},\dots,\widehat{i}_{T}\}$. So, the regret becomes:
$R_{\textsc{DES}}(T)\leq\textsf{OPT}-\textsf{Explr}-\widehat{\textsf{DP}}+\widehat{\textsf{DP}}-\textsf{DP}$
The reason why we needed the “stepping stone” of $\widehat{\textsf{DP}}$ in the regret evaluation is that the total reward our algorithm picks up is expressed in terms of $(r_{i},b_{i})$ but the dynamic programming guarantees are expressed in terms of $(\widehat{r}_{i},\widehat{b}_{i})$.
We next define events $\mathcal{E}_{r,i}=\{|\widehat{r}_{i}-r_{i}|\leq\delta\}$ and $\mathcal{E}_{b,i}=\{|\widehat{b}_{i}-b_{i}|\leq\delta\}$ for all $i\in[K]$. Then, conditional on the event =$\mathcal{E}=\{\cap_{i\in[K]}(\mathcal{E}_{r,i}\,\,\text{and}\,\,\mathcal{E}_{v,i})\}$ and due to Lemma 3.3 and Lemma 3.4, the regret picked up for all the remaining rounds after feeding estimates $\{(\widehat{r}_{i},\widehat{b}_{i})\}_{i\in[K]}$ to the dynamic programming procedure is at most $2\delta T$. As a result, from the law of total probability, the regret for all $T$ rounds: